Which of the ordered pairs
(3, 1), (0, –4), (–4, 0), (–3, –7)are solutions for the equation x – y = 4?

Answer by ikleyn(52787)


The pair  (x,y) = (3,1)  gives   x - y = 3-1 = 2.

The pair  (x,y) = (0,-4)  gives   x - y = 0-(-4) = 4.

The pair  (x,y) = (-4,0)  gives   x - y = -4-0 = -4.

The pair  (x,y) = (-3,-7)  gives   x - y = -3-(-7) = 7-3 = 4.

Answer.  Two pairs  (x,y) = (0,-4)  and  (x,y) = (-3,-7)  give   x - y = 4.